If you double the radius of an ordinary (non-fractal) solid rubber ball, the mass of the ball increases how many times?

Question 2 options:

2

4

8

16

1/2
Answer #1

we know that mass is given by:

Mass = density*Volume

Since density of rubber ball will remains constant, So

Volume of solid rubber ball = 4*pi*r^3/3

Mass = density*(4*pi*r^3/3)

Which means mass is directly proportional to the cube of radius of ball

M2/M1 = (r2/r1)^3

given that r2 = 2*r1

So,

M2 = M1*(2*r1/r1)^3

M2 = 8*M1

So mass will be multiplied by a factor of 8.

Correct option is C.

Please Upvote.